# PDDFormer: Pairwise Distance Distribution Graph Transformer for Crystal Material Property Prediction

### v4

@@ 4.6 Network architecture
CFID (2018) 14 0.30 240 220 0.53
CGCNN 63 0.20 78 170 0.41
SchNet 45 0.19 47 140 0.43
MEGNET 47 0.145 58 84 0.34
GATGNN 47 0.17 56 120 0.51
ALIGNN 33.1 0.142 37 76 0.31
M3GNet 39.0 0.145 41 95 0.36
Matformer 32.5 0.137 35 64 0.30
PotNet 29.4 0.127 32 55 0.27
CrysMMNet 28.0 0.128 34 0.278
CrysDiff (2024) 29.0 0.131 34 62 0.287
Crystalformer 30.6 0.128 32 46 0.274
eComFormer 28.4 0.124 32 44 0.28
iComFormer 27.2 0.122 28.8 47 0.26
UPDDFormer 27.6 0.127 29.4 35.6 0.269
WPDDFormer 26.9 0.120 28.2 32.6 0.251
Table 1: Comparison between Zeoformer and other baselines in terms of test MAE on JARVIS dataset. The best results are shown in bold and the second best results are shown with underlines.
Formation Energy Band Gap Bulk Moduli Shear Moduli
Method eV/atom eV log(GPa) log(GPa)
CGCNN (2018) 31 0.292 0.047 0.077
SchNet (2018) 33 0.345 0.066 0.099
MEGNET (2019) 30 0.307 0.060 0.099
GATGNN (2020) 33 0.280 0.045 0.075
ALIGNN (2021) 22 0.218 0.051 0.078
M3GNet (2022) 24 0.247 0.050 0.087
Matformer (2022) 21.0 0.211 0.043 0.073
PotNet (2023) 18.8 0.204 0.040 0.065
CrysMMNet (2023) 20.0 0.197 0.038 0.062
Crystalformer (2024) 18.6 0.198 0.0377 0.0689
eComFormer (2024) 18.16 0.202 0.0417 0.0729
iComFormer (2024) 18.26 0.193 0.0380 0.0637
UPDDFormer 18.31 0.196 0.0393 0.0686
WPDDFormer 16.61 0.189 0.0336 0.0617
Table 2: Comparison between Zeoformer and other baselines in terms of test MAE on MP dataset.
Node-wise transformer block. Building upon the constructed graph, we aggregate the node information. Let $a_{i}^{l}$ represent the input feature vector of node $i$ at layer $l$ in PDDFormer. The information propagation scheme for layer $l$ is described as follows. We use $\oplus$ and $\odot$ to denote concatenation and element-wise product. The message from node $j$ to $i$ is formed by the corresponding query $q_{ij}^{l}$, key $k_{ij}^{l}$, and value features $v_{ij}^{l}$ as follows :
$\begin{array}[]{@{}c@{}c@{}c@{}}q_{ij}^{l}=LQ\left(a_{i}^{l}\right),\>k_{ij}^{l}=\left(LK\left(a_{i}^{l}\right)\oplus LK\left(a_{j}^{l}\right)\oplus LE\left(e_{ij}^{h}\right)\right),\\ v=\left(LV\left(a_{i}^{l}\right)\oplus LV\left(a_{j}^{l}\right)\oplus LE\left(e_{ij}^{h}\right)\right),\\ v_{ij}^{l}=v\odot sigmoid\left(LN_{orm}\left(LN_{att}\left(v\right)\right)\right),\\ att^{l}=\frac{q_{ij}^{l}\odot LN_{K}\left(k_{ij}^{l}\right)}{\sqrt{d_{q_{ij}^{l}}}},\\ m_{ij}^{h}=q_{ij}^{l}+sigmoid\left(BN\left(att^{l}\right)\right)\odot LN_{V}\left(v_{ij}^{l}\right)\end{array}$ (5)
where LQ, LK, LV, and LE are the linear transformations for query, key, value, and edge features. $LN_{K}$, $LN_{V}$ are the nonlinear transformations for key and value, including two linear layers and an activation layer in between. $LN_{update}$ represents the linear transformation for updating messages, and $LNorm$ denotes the layer normalization Ba (2016) operation. BN denotes the batch normalization layer Ioffe (2015), and $d_{q_{ij}^{l}}$ is the dimension of $q_{ij}^{l}$.
$\begin{array}[]{@{}c@{}c@{}}M_{i}^{l}=\sum_{j\in A_{i}}\sum_{h}LN_{sum}\left(m_{ij}^{h}\right),\\ A_{i}^{l+1}=softplus\left(a_{i}^{l}+BN\left(M_{i}^{l}\right)\right)\end{array}$ (6)
where $LN_{msg}$ is the linear transformation used for updating the edge messages.
PDD message passing block. $\mathcal{A}^{l}$ and $PDD^{l}$ represent the atomic features and 3D periodic pattern encoding at layer $l$, respectively. Its message-passing mechanism is as follows:
$\begin{array}[]{@{}c@{}c@{}c@{}}\mathcal{PDD}^{l+1}=\mathcal{PDD}^{l}+\mathcal{A}^{l+1},\\ A_{1},A_{2}=LN_{PDD}\left(BN\left(PDD^{l+1}\right)\right),\\ A^{l+1}=A^{l}+LN_{A2}\left(LN_{A1}\left(A_{1}\right)\odot Drop\left(GELU\left(A_{2}\right)\right)\right)\end{array}$ (7)
In this process, we update $A^{l+1}$ using residual connections He et al. (2016).
Finally, we use average pooling to aggregate the features of all nodes in the graph, followed by a nonlinear layer, and then a linear layer to obtain the scalar output of the graph as described above. A detailed description of the PDDFormer architecture can be found in Appendix C.
5 Experiments
@@
We conducted experiments on two material benchmark datasets, namely the Materials Project Chen et al. (2019) and Jarvis Choudhary et al. (2020) datasets. Detailed descriptions of the datasets can be found in Appendix A. More information about the experimental settings of PDDFormer can be found in Appendix D. Baseline methods include CFID Choudhary et al. (2018), CGCNN Xie and Grossman (2018), SchNet Schütt et al. (2017), MEGNET Chen et al. (2019), GATGNN Louis et al. (2020), ALIGNN Choudhary and DeCost (2021), M3GNet Chen and Ong (2022), Matformer Yan et al. (2022), PotNet Lin et al. (2023), CrysMMNet Das et al. (2023), CrysDiff Song et al. (2024), Crystalformer Taniai et al. (2024), and ComFormer Yan et al. (2024a). For all baselines on the material datasets, we report the results provided in the cited papers.
5.1 Experimental results
JARVIS.
The quantitative results for JARVIS Choudhary et al. (2020) are shown in Table LABEL:sample-table1. WPDDformer achieves the best performance across all tasks. Notably, WPDDFormer and UPDDFormer outperform eComFormer by 26% and 19% respectively in the Ehull task.
The Materials Project (MP).
The experimental results on MP Chen et al. (2019) are shown in Table LABEL:sample-table2. WPDDformer performs significantly better than previous works across all tasks, with a $10.8$% improvement over the second-best model in the bulk moduli task. Additionally, the excellent prediction accuracy of WPDDFormer in the bulk modulus and shear modulus tasks, using only $4,664$ training samples, demonstrates the expressiveness and robustness of WPDD multi-edge crystal graphs under limited training samples.
Table 3: Efficiency comparison with ConFormer on the Jarvis Formation Energy task. We show the training time per epoch, total training time, time complexity, GPU memory consumption, and total number of parameters. The experiments were conducted using a 3090 RTX 24GB GPU.
Models Time/epoch Total GPU memory usage Complexity Model Para.
eConformer 120s 16.7h 18GB $O(nk)$ 12.4M
iConformer 129s 25.0h 12GB $O(nk)$ 5.0M
WPDDFormer 98s 10.9h 8.5GB $O(nk)$ 6.76M
Method Num. Block Ehull Bulk
NO PDD Block 4,0 39.2 0.0410
without PDD 4,3 36.3 0.0400
UPDDFormer 3,2 37.4 0.0446
UPDDFormer 4,3 35.6 0.0392
WPDDFormer 3,2 34.0 0.0336
WPDDFormer 4,3 32.6 0.0341
Table 4: Num. Block represents the number of Node-wise transformer blocks and PDD message passing blocks.
Overall, our methods are compared with $14$ existing methods across the two datasets. Our WPDDFormer consistently outperforms all methods in all tasks. Additionally, WPDDFormer shows a significant improvement in prediction accuracy compared to UPPDFormer. This improvement is not only because the WPDD graph structure is complete and continuous, while UPPD can only ensure continuity, but also because UPPD requires dimensional alignment as mentioned in 4.6, which results in some loss of the expression of global information about the unit cell.
Efficiency This experiment reports the training and inference times for WPDDFormer and ConFormer using the best model configurations. We also report the total number of parameters for each model. As shown in Table 3, all these models have a time complexity of $O(nk)$, where $n$ represents the number of atoms in the unit cell and $k$ represents the average number of neighbors. The data in the table is averaged over three experiments. Although WPDDFormer has a higher parameter count compared to iConFormer, its training time overhead is significantly lower than that model, and it uses less GPU memory. Its memory usage is only $70.8$% of iConFormer and $47.2$% of eConFormer. This demonstrates that our WPDDFormer achieved significantly superior experimental results with lower computational cost and faster computation speed. Additional four tasks from the JARVIS dataset are documented in Appendix E.
5.2 Ablation studies
In this section, we demonstrate the impact of introducing (W/U)PDD on the representation learning of crystal materials through ablation studies. Specifically, we conducted experiments on the MP and JARVIS datasets, using testing mean absolute error (MAE) as the quantitative evaluation metric, comparing the results for Bulk Moduli and Ehull tasks, as shown in Table 4.
By comparing (W/U)PDDFormer models with different numbers of Node-wise Transformer Blocks and PDD Message Passing Blocks to models without (W/U)PDD information but retaining the $PDD$ message passing blocks, we validate the importance of (W/U)PDD. The results show that compared to models without the $PDD$ message passing blocks, WPDDForemer achieved improvements of $18.0\%$ and $16.8\%$ in the Bulk Moduli and Ehull tasks, respectively. Compared to models that retain only the $PDD$ message passing blocks but lack (W/U)PDD information, we achieved improvements of $16.0\%$ and $10.2\%$ in these two tasks, respectively.
6 Conclusion and future work
In this study, we integrated WPDD and UPDD into the representation of crystal structures, achieving a complete and continuous construction of crystal graphs. This resolves the ambiguity in crystal graph representations for predicting the properties of crystalline materials and bridges the gap between traditional crystal descriptors and dynamic atomic behavior. Experimental results demonstrate the significant advantage of our WPDDFormer in various property prediction tasks. Ensuring the completeness and continuity of crystal graphs after incorporating angular information is a problem that will be further explored in the future.

### v5

@@ D.2 Proofs of invariance
No PDD 27.0 0.122 28.0 42.6 0.261
WPDDFormer 25.7 0.119 27.6 35.5 0.249
(a) JARVIS-FE
(b) JARVIS-Ehull
(c) MP-FE
(d) MP-Bulk
Figure 4: The coefficient of determination for WPDDFormer’s predictions is presented. The scatter plots reflect the differences between our predicted values and the actual values, while the bar charts show the frequency of the values.
Figure 5: The different neighbor selection under slight perturbations.
Appendix E More experimental
E.1 Complete experimental data.
In this section, we provide the complete baseline model data and additional experimental results compared in the main text.
@@ Appendix E More experimental
Baseline methods include CFID  choudhary2018machine, CGCNN  xie2018crystal, SchNet  schutt2017schnet, MEGNET  chen2019graph, GATGNN  louis2020graph, ALIGNN  choudhary2021atomistic, M3GNet  chen2022universal, Matformer  yan2022periodic, PotNet  lin2023efficient, CrysMMNet  das2023crysmmnet, CrysDiff  song2024diffusion, Crystalformer  taniai2024crystalformer, and ComFormer  yan2024complete. For all baseline methods on the material datasets, we report the results provided in the cited papers or those reproduced based on their official codes.
JARVIS and MP.
As shown in Tables 7 and 8, WPDDFormer achieved the best results across all 13 tasks on these two datasets, a performance that no previous model has achieved, highlighting the strong superiority of WPDDFormer. For the Bulk Moduli (Kv) and Shear Moduli (Gv) tasks in JARVIS, 19,680 training samples were used, while 9,066 and 11,375 training samples were used for the SLME (%) and Spillage tasks, respectively. For the Bulk Moduli (Kv) and Shear Moduli (Gv) tasks in MP, 5,450 training samples were used. This demonstrates WPDDFormer’s adaptability to tasks with varying data scales. UPDDFormer surpasses WPDDFormer in inference speed while also demonstrating strong performance compared to previous models, achieving second-best results in 10 out of the 13 tasks.
We selected $R_{2}$ (coefficient of determination) as the evaluation metric. $R_{2}$ is one of the indicators for measuring the goodness-of-fit of a model, representing the extent to which the predicted values explain the actual values. As shown in Figure 4, we conducted experiments on the formation energy and Ehull metrics from the JARVIS dataset, as well as the formation energy and bulk moduli from the MP dataset. The experimental results show that the $R_{2}$ accuracy of our model has exceeded 0.99 in datasets with a larger number of training elements. Even on the Bulk property with smaller data volumes, the model performs well, demonstrating its superior performance.
E.2 The number of neighbors ( k ) of WPDD
In this section, we investigate the effect of the number of neighbors with different cutoff radii on the WPDD experiment. In the main text, we set the number of neighbors $k=92$, which matches the dimensionality of the atomic feature embeddings used by CGCNN. However, in practical applications, it is necessary to determine a sufficiently large $k$ in advance to ensure completeness for any test crystal, especially in extreme cases where $k$ must be greater than the number of atoms in any test crystal  yan2024complete. Therefore, to ensure completeness on the JARVIS and MP datasets, we calculated the maximum number of atoms in each dataset, which is $140$ for JARVIS, $152$ for MP’s bulk and shear, and $296$ for the rest. As a result, in the experiments, the maximum number of neighbors for the JARVIS and MP datasets with different cutoff radii were selected as $150$, $160$, and $300$, respectively. Therefore, we can ensure completeness for general-purpose datasets.
From the experimental results in Tables 9 and 10, it can be seen that the choice of neighbors with different cutoff radii causes fluctuations in the model’s performance. However, its performance still shows the best results compared to the other models presented in the main text.
E.3 More ablation
To verify the effectiveness of incorporating PDD descriptors into crystal graph construction, we present in Table 11 the impact on experimental results on the JARVIS dataset. With minimal changes to the WPDDFormer model parameters, we investigated the effect of removing WPDD information on the experiments. It can be observed that after incorporating WPDD information, the model’s performance significantly improved, highlighting the importance of introducing WPDD information.
Appendix F Continuous tolerance $\mathcal{T}$
Given that the experimentally measured unit cell and atomic coordinates are inevitably affected by atomic vibrations and measurement noise, slight perturbations in the atomic coordinates may occur. Therefore, during the construction of the multi-edge crystal graph, when selecting the t nearest edges within a cutoff radius, the chosen neighboring nodes may change, as shown in Figure 5 (a) and (b). This variation in neighboring nodes $j$ leads to changes in the atomic information of the neighboring nodes, resulting in discontinuities in the construction of the multi-edge crystal graph. To eliminate the influence of atomic perturbations on the neighbor selection and ensure the continuity of the crystal graph under perturbations, we define the concept of continuous tolerance $\mathcal{T}$ to guarantee the continuity of the constructed crystal graph.
Since the distances of atomic perturbations are typically on the order of sub-angstrom (Å), specifically, in common atomic structures or crystals, slight perturbations are generally less than $\mathcal{T}<10^{-2}$. For larger perturbations (e.g., $\mathcal{T}=10^{-2})$, the continuity issue may no longer be effective. Therefore, when selecting neighbors, we can set a continuous tolerance value $\mathcal{T}$ in advance. When we select the t nearest edges within the cutoff radius, if the distance of the (t+1)-th edge from node $i$ minus the distance of the t-th edge is less than the continuous tolerance, i.e., $d_{k+1}-d_{k}<T$, we include this neighbor in the graph construction as well. This process continues until the distance of the (n+1)-th edge from node i minus the distance of the n-th edge exceeds the continuous tolerance cutoff, i.e., $d_{n+1}-d_{n}>\mathcal{T}$. The n neighbors at this point are the selected nodes for constructing the crystal graph, as shown in Figure 5 (d), resulting in the final neighbor selection shown in Figure 5 (c). This approach ensures that the neighbor selection in the crystal graph construction does not change under atomic perturbations, and by using Lemma 1, we prove the continuity of the (W/U)PDD crystal graph we have constructed.
Appendix G Limitation
PDD demonstrates exceptional discriminative capability: with the number of neighbors k=100, it can distinguish all known crystal structures in the Cambridge Structural Database (CSD). However, a theoretical scenario exists where, with an infinitely large unit cell, no matter the value of k, there will always be a larger unit cell that renders k insufficient to ensure completeness. Therefore, absolute completeness cannot be achieved in a theoretical sense. Nonetheless, such extreme scenarios exist only in theoretical derivations and are exceedingly improbable, even in theory. More importantly, these situations do not occur in the real world. Thus, our approach satisfies the general completeness requirements for almost all materials under perturbations.